how much kinetic energy does a 2500kg minivan traveling at 35m/s have?


1
Expert's answer
2021-01-10T18:25:18-0500

By the definition of the kinetic energy, we have:


"KE=\dfrac{1}{2}mv^2=\dfrac{1}{2}\cdot 2500\ kg\cdot(35\ \dfrac{m}{s})^2=1.53\cdot10^6\ J."

Answer:

"KE=1.53\cdot10^6\ J."